OpenStudy (anonymous):

Simplify. 4√2 + 7√2 - 3√2

OpenStudy (anonymous):

Is that square root of 2?

OpenStudy (midhun.madhu1987):

It is as simple as 4x + 7x - 3x

OpenStudy (anonymous):

Note that \[ 4\sqrt{2} + 7\sqrt{2} - 3\sqrt{2} = (4 + 7 - 3) \sqrt{2}, \]because you can distribute multiplication. Can you finish the rest

OpenStudy (anonymous):

How do you do it

OpenStudy (anonymous):

Oh okay, yeah i can